Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
mantovani, nardecchia, ruggi, magazzu - 17:19 Thursday 21 September 2023 (61741) Print this report
DAS differential tuning - PART 2

This morning, we proceeded with the DAS differential tuning. In order to try to evaluate the effect of each DAS ring, the NI outer ring has been decreased by 15 %. 

08.17 UTC: Locked at LN2

08.48 UTC: NI Outer ring power change (-15 % wrt the STEP_1S ) 

 

STEP2_S

 

08.48 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.068

0.078*(1+15/100)=0.090

1.615*(1+15/100)=1.857

on the pickoff

0.418

0.0127*(1+15/100)= 0.0146

0.263*(1+15/100)=0.302

N

on the ITF

0.091

0.219*(1-15/100)+0.08=

0.266 

1.172*(1-15/100)+1.8 =2.8

on the pickoff

0.561

0.036*(1-15/100)+0.013=

0.044

0.191*(1-15/100)+0.293=
0.455

09.13 UTC- 9.18 UTC: Manuel tested a new filter

09.18 UTC: ITF unlocked

09.25- 09.46 : Diego P. went to the CB on the platform.

At the next relock, the ITF working point appeared degraded (see fig.1).

Thus, in order to investigate the ratio between NI inner and outer ring, the NI Inner ring power has also been decreased by the same amount (-15 %) .

10.40 UTC:  NI Inner ring power change (-15 % wrt the STEP_2S ). 

 

STEP3_S

 

10.40 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.068

0.078*(1+15/100)=0.090

1.615*(1+15/100)=1.857

on the pickoff

0.418

0.0127*(1+15/100)= 0.0146

0.263*(1+15/100)=0.302

N

on the ITF

0.091

0.1861*(1-15/100)+0.08=0.238

1.172*(1-15/100)+1.8 =2.8

on the pickoff

0.561

0.026+0.013=0.039

0.191*(1-15/100)+0.293=
0.455

Also this step had a negative impact on the ITF WP, B1p was very high and we were not able to lock in low noise 2 due to unbalance of B1P PC 56 MHz.

Thus, with the ITF in CARM NULL 1F, the NI DAS inner and outer ring powers have been increased again, at 12.08 UTC.

 

STEP4_S

(= STEP1_S)

 

12.08 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.068

0.078*(1+15/100)=0.090

1.615*(1+15/100)=1.857

on the pickoff

0.418

0.0127*(1+15/100)= 0.0146

0.263*(1+15/100)=0.302

N

on the ITF

0.091

0.219*(1-15/100)+0.08=

0.266 

1.379*(1-15/100)+1.8 =2.972

on the pickoff

0.561

0.036*(1-15/100)+0.013=

0.044

0.225*(1-15/100)+0.293=

0.484

13.15 UTC: unlocked

13.24:WI DAS powers have been increased by 15%.

 

STEP5_S

 

12.39 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.068

0.090*(1+15/100)=0.104

1.615*(1+15/100)=2.13

on the pickoff

0.418

0.0146*(1+15/100)= 0.0169

0.302*(1+15/100)=0.347

N

on the ITF

0.091

0.219*(1-15/100)+0.08=

0.266 

1.379*(1-15/100)+1.8 =2.972

on the pickoff

0.561

0.036*(1-15/100)+0.013=

0.044

0.225*(1-15/100)+0.293=

0.484

At 14.05 UTC, the ITF locked in LN2.

It seems to be a good WP (see fig.2) but it is needed to dedicate some time to confirm it.   

We left these DAS powers identified as STEP 5_S for the afternoon shift. 

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×